# RF-DETR: Neural Architecture Search for Real-Time Detection Transformers

### v1

@@ 5 Conclusion
In this paper, we introduce RF-DETR, a state-of-the-art NAS-based method for fine-tuning specialist end-to-end object detectors for target datasets and hardware platforms. Our approach outperforms prior state-of-the-art real-time methods on COCO and RF100-VL, improving upon D-FINE (nano) by 5% AP on COCO. Moreover, we highlight that current architectures, learning rate schedulers and augmentation schedulers are tailored to maximize performance on COCO, suggesting that the community should benchmark models on diverse, large-scale datasets to prevent implicit overfitting. Lastly, we highlight the high variance in latency benchmarking due to power throttling and propose a standardized protocol to improve reproducibility.
@@ Appendix A Implementation Details
Training Hyperparamters. RF-DETR extends LW-DETR (chen2024lw) for Neural Architecture Search. We highlight key differences in our training procedure below. First, we pseudo-label Objects365 (objects365) with SAM2 (ravi2024sam2) to allow us to pre-train the segmentation and detection heads on the same data. We use a learning rate of 1e-4 (LW-DETR uses 4e-4), and a batch size of 128 (LW-DETR uses the same). Similar to DINOv3 (simeoni2025dinov3), we use an EMA scheduler since this is necessary for EMA’s proper function. However, unlike DINOv3, we omit learning-rate warm-up. We clip all gradients greater than 0.1 and apply a per-layer multiplicative decay of 0.8 to preserve information (especially the earlier layers) in the DINOv2 backbone. We place our window attention blocks between layers {0, 1, 3, 4, 6, 7, 9, 10}, while LW-DETR places their window attention blocks between layers {0, 1, 3, 6, 7, 9}. Although we have the same number of windows, contiguous windowed blocks don’t require an additional reshape operation, making our implementation slightly more efficient. Further, we train with more multi-scale resolutions (0.5 to 1.5 scale) than LW-DETR (0.7 to 1.4 scale) to ensure that the augmentation is symmetric around the default scale. Notably, we add resolution as a “tunable knob” in our NAS search space, while LW-DETR uses it as a form of data augmentation. Our model training and inference code is available on GitHub.
@@
Pareto-Optimal Model Configurations on COCO. We present the Pareto-Optimal RF-DETR and RF-DETR-Seg configs in Tables 7 and 8. We highlight notable trends about RF-DETR’s Pareto-Optimal architectures in Appendix H.
@@ Appendix B Ablation on Query Tokens and Decoder Layers
We train RF-DETR (nano) with 300 object queries, following standard practice for real-time DETR-based object detectors. However, many datasets contain fewer than 300 objects per image. Therefore, processing all 300 queries can be computationally wasteful. LW-DETR (tiny) demonstrates that training with fewer queries can improve the latency-accuracy tradeoff. Rather than deciding on the optimal number of queries apriori, we find that we can drop queries at test time without retraining by discarding the lowest-confidence queries ordered by the confidence of the corresponding token at the output of the encoder. As shown in Figure 4, this yields meaningful latency-accuracy tradeoffs. In addition, prior work (zhao2024rtdetr) demonstrates that decoder layers can be pruned at test time, since each layer is supervised independently during training. We find that it is possible to remove all decoder layers, relying solely on the initial query proposals from the two-stage DETR pipeline. In this case, there is no cross-attention to the encoder states or self-attention between queries, leading to a substantial runtime reduction. The resulting model resembles a single-stage YOLO-style architecture without NMS. As shown in Figure 4, eliminating the final decoder layer reduces latency by 10% with only a 2 mAP drop in performance.
@@ Appendix C Benchmarking FLOPs
We benchmark FLOPs for RF-DETR  GroundingDINO, and YOLO-E with PyTorch’s FlopCounterMode. We find that FlopCounterMode closely reproduces FLOPs counts obtained with custom benchmarking tools for YOLOv11, D-FINE, and LW-DETR. In practice, we also find that it provides more reliable results than CalFLOPs (calflops). Notably, LW-DETR’s FLOPs count is roughly twice that of the originally reported result (cf. Table 9). We posit that this discrepancy can be attributed to LW-DETR reporting MACs instead of FLOPs. We rely on the officially reported FLOPs counts from YOLOv11, YOLOv8, D-FINE, and RT-DETR.
@@ Appendix D Impact of Class-Names on Open-Vocabulary Detectors
We evaluate the impact of fine-tuning open-vocabulary detectors like GroundingDINO with class names on RF100-VL in Table 10. Intuitively, GroundingDINO’s vision-language pre-training is more useful when we prompt with class names (e.g. car, truck, bus) instead of class indices (e.g. 0, 1, 2). Using class names at finetune time therefore provides more information to the VLM about the underlying data than is available to non-VLM detectors, potentially leading to better downstream performance. However, we find that fine-tuning GroundingDINO on RF100-VL yields nearly identical performance in both cases, suggesting that naively fine-tuning the end-to-end model mitigates the benefits of open-vocabulary pre-training. Future should investigate ways of effectively fine-tuning VLMs to preserve foundational pre-training.
@@
GroundingDINO (liu2023grounding) w/ Standard Class Names T 173.0M 1008.3 309.9 62.3 88.8 67.8 39.2 57.7 69.5
GroundingDINO (liu2023grounding) w/ Class Index Names T 173.0M 1008.3 309.9 62.5 88.2 68.3 40.0 58.4 70.3
@@ Appendix E Benchmarking Larger Model Variants
Detectors like LW-DETR (chen2024lw) and D-FINE (peng2024dfine) hand-design larger variants to scale up a model family. In contrast, NAS-based architectures like RF-DETR automatically discover scaling strategies through grid-based search. We analyze two families of RF-DETR models derived from distinct scaling strategies: one based on a DINOv2-S backbone and another based on a DINOv2-B backbone. To evaluate how well each family scales, we compare their NAS-generated Pareto curves against those of D-FINE. Specifically, at each D-FINE size, we identify the RF-DETR variant with the same backbone that maximizes performance at a comparable latency. For example, when comparing to D-FINE (small), we select the RF-DETR model that offers the best accuracy without exceeding D-FINE (small)’s latency.
As shown in Table 11, the DINOv2-S backbone family initially surpasses D-FINE in mAP@50:95 but fails to maintain this advantage at larger model sizes, suggesting that its scaling strategy is less effective than D-FINE’s manual design. In contrast, the DINOv2-B backbone family shows the opposite trend, where the performance gap between D-FINE and RF-DETR narrows as latency increases. This implies that at higher latencies, the DINOv2-B based RF-DETR models could surpass D-FINE (and indeed RF-DETR (2x-large) outperforms D-FINE on mAP 50:95). Importantly, expanding the D-FINE model family would require substantial additional engineering effort, whereas extending the RF-DETR model family is straightforward; higher-latency variants can be sampled directly from the same NAS search without re-training. We present the COCO and RF100-VL of our larger variants in Tables 12, 13, and 14. We also include an RF-DETR Max variant on each dataset to show our method’s maximum performance with latency less than 100ms, a scale other model families don’t reach.
@@
YOLOv8 (yolov8) L 43.7M 165.2 8.0 49.5 64.7 54.0 30.2 55.1 68.5
YOLOv11 (yolov11) L 25.3M 86.9 6.5 49.9 64.9 54.5 30.4 55.9 68.1
YOLOv8 (yolov8) XL 68.2M 257.8 11.3 50.5 65.6 55.1 30.0 56.2 69.5
YOLOv11 (yolov11) XL 56.9M 194.9 10.5 50.9 66.1 55.4 31.5 56.6 68.7
@@
RT-DETR (zhao2024rtdetr) R50 42M 136 8.5 55.0 73.3 59.8 37.9 59.7 71.6
LW-DETR (chen2024lw) L 46.8M 137.5 6.9 56.1 74.6 61.0 37.1 60.4 73.0
D-FINE (peng2024dfine) L 31M 91 7.5 57.2 74.9 62.2 40.6 61.4 73.7
@@
RT-DETR (zhao2024rtdetr) R101 76M 259 12.0 56.1 74.5 61.1 38.1 60.4 73.4
LW-DETR (chen2024lw) XL 118.0M 342.5 13.0 58.3 76.9 63.3 40.2 63.3 74.7
D-FINE (peng2024dfine) XL 62M 202 11.5 59.3 76.8 64.6 42.1 64.2 76.3
@@
YOLOv8 (yolov8) L 46.0M 220.5 9.7 39.0 60.5 41.7 18.0 44.7 57.8
YOLOv11 (yolov11) L 27.6M 132.2 8.3 39.5 61.5 42.1 18.6 45.5 59.4
YOLOv8 (yolov8) XL 71.8M 344.1 14.0 39.5 61.3 42.1 18.9 45.6 58.8
YOLOv11 (yolov11) XL 62.1M 296.4 13.7 40.1 62.4 42.6 18.8 46.4 60.1
@@
YOLOv8 (yolov8) L 43.7M 165.2 7.9 56.5 82.1 61.1 7.1 46.0 48.9
YOLOv11 (yolov11) L 25.3M 86.9 6.4 56.5 82.2 61.0 6.4 45.5 49.0
YOLOv8 (yolov8) XL 68.2M 257.8 11.2 56.5 82.3 61.0 6.6 45.7 47.9
YOLOv11 (yolov11) XL 56.9M 194.9 10.3 56.2 81.7 60.8 6.1 45.9 48.1
@@
RT-DETR (zhao2024rtdetr) R50 42M 136 8.4 61.7 87.7 66.9 38.1 57.1 69.4
LW-DETR (chen2024lw) L 46.8M 137.5 6.8 61.5 87.4 67.0 37.1 56.4 69.0
D-FINE (peng2024dfine) L 31M 91 7.5 61.6 86.4 67.2 37.8 56.5 70.1
RF-DETR (Ours) L 34.1M 119.1 6.3 62.3 88.2 68.0 36.4 57.3 70.6
RF-DETR (Ours) w/ Fine-Tuning L 34.1M 119.1 6.3 62.6 88.4 68.2 37.0 57.5 70.5
RT-DETR (zhao2024rtdetr) R101 76M 259 11.9 61.0 87.4 66.2 36.6 56.3 68.2
LW-DETR (chen2024lw) XL 118.0M 342.5 13.0 62.1 87.9 67.6 37.4 57.1 70.2
D-FINE (peng2024dfine) XL 59.3 76.8 11.4 62.2 86.9 68.0 37.6 57.4 69.7
RF-DETR (Ours) XL 35.0M 199.0 9.8 62.7 88.5 68.5 39.3 58.4 70.4
RF-DETR (Ours) w/ Fine-Tuning XL 35.0M 199.0 9.8 63.1 88.6 69.0 39.6 58.5 70.8
@@
Appendix F Impact on NAS Fine-Tuning on COCO
@@
Appendix G Impact of Fixed Architecture on RF100-VL
We evaluate the impact of transferring a NAS architecture optimized for COCO to RF100-VL in Table 17. We find that these fixed architecture models perform remarkably well without further dataset-specific NAS. Specifically, RF-DETR (large) model with a fixed architecture achieves the best performance among all prior real-time models on COCO. However, dataset-specific NAS yields significant additional gains. Notably, the improvement from LW-DETR to the fixed architecture is comparable to the improvement from the fixed architecture to the NAS-optimized model on the target dataset for N, S, and M scale models.

### v2

@@ 5 Conclusion
In this paper, we introduce RF-DETR, a state-of-the-art NAS-based method for fine-tuning specialist end-to-end object detectors for diverse target datasets and hardware platforms. Our approach outperforms prior state-of-the-art real-time methods on COCO and RF100-VL, improving upon D-FINE (nano) by 5% AP on COCO. Moreover, we highlight that current architectures, learning rate schedulers and augmentation schedulers are tailored to maximize performance on COCO, suggesting that the community should benchmark models on diverse, large-scale datasets to prevent implicit overfitting. Lastly, we highlight the high variance in latency benchmarking due to power throttling and propose a standardized protocol to improve reproducibility.
Acknowledgements
This work was supported in part by compute provided by NVIDIA DGX. We’d like to thank Brad Dwyer for reviewing early versions of our manuscript.
@@ Appendix A Implementation Details
Training Hyperparamters. RF-DETR extends LW-DETR (Chen et al., 2024a) for Neural Architecture Search. We highlight key differences in our training procedure below. First, we pseudo-label Objects-365 (Shao et al., 2019) with SAM2 (Ravi et al., 2024) to allow us to pre-train the segmentation and detection heads on the same data. We use a learning rate of 1e-4 (LW-DETR uses 4e-4), and a batch size of 128 (LW-DETR uses the same). Similar to DINOv3 (Siméoni et al., 2025), we use an EMA scheduler since this is necessary for EMA to function. However, unlike DINOv3, we omit learning-rate warm-up. We clip all gradients greater than 0.1 and apply a per-layer multiplicative decay of 0.8 to preserve information (especially the earlier layers) in the DINOv2 backbone. We place our window attention blocks between layers {0, 1, 3, 4, 6, 7, 9, 10}, while LW-DETR places their window attention blocks between layers {0, 1, 3, 6, 7, 9}. Although we have the same number of windows, contiguous windowed blocks don’t require an additional reshape operation, making our implementation slightly more efficient. Further, we train with more multi-scale resolutions (0.5 to 1.5 scale) than LW-DETR (0.7 to 1.4 scale) to ensure that the augmentation is symmetric around the default scale. Notably, we add resolution as a “tunable knob” in our NAS search space, while LW-DETR uses it as a form of data augmentation. Our model training and inference code is available on GitHub.
@@
Pareto-Optimal Model Configurations in COCO. We present the Pareto-Optimal RF-DETR and RF-DETR-Seg configurations in Tables 7 and 8. We highlight notable trends about RF-DETR’s Pareto-Optimal architectures in Appendix L.
@@
Parameter Sampling Grid. Lastly, we present our sampling grid for training and inference below. Importantly, we only drop decoder layers and queries during inference. We uniformly sample configurations during training and perform grid search over all configurations during inference to find Pareto-Optimal model configurations. RF-DETR’s total training time is roughly two to four times as long as a non-NAS baseline, depending on the target dataset. However, RF-DETR can generate all size configurations from this single training run, while other non-NAS baselines must be re-trained for each new model size. We evaluate 6,468 network configurations (11 resolutions * 7 patch sizes * 7 decoder layers * 3 windows * 4 query settings) during architecture search. We estimate that this search takes approximately 10,000 GPU Hours (200 T4 GPUs * 48 hours).
Training Configurations
  • Image Resolutions: 320, 384, 448, 512, 576, 640, 704, 768, 832, 896, 960
  • Patch Sizes: 8, 10, 12, 16, 20, 24, 32
  • Number of Decoder Layers: 6
  • Number of Windows: 1, 2, 4
  • Number of Queries: 300
Inference Configurations
  • Image Resolutions: 320, 384, 448, 512, 576, 640, 704, 768, 832, 896, 960
  • Patch Sizes: 8, 10, 12 ,14, 16, 20, 24, 32
  • Number of Decoder Layers: 0, 1, 2, 3, 4, 5, 6
  • Number of Windows: 1, 2, 4
  • Number of Queries: 50, 100, 200, 300
@@ Appendix B Ablation on Query Tokens and Decoder Layers
We train RF-DETR (nano) with 300 object queries, following standard practices for real-time DETR-based object detectors. However, many datasets contain fewer than 300 objects per image. Therefore, processing all 300 queries can be computationally wasteful. LW-DETR (tiny) demonstrates that training with fewer queries can improve the latency-accuracy tradeoff. Rather than deciding on the optimal number of queries apriori, we find that we can drop queries at test time without retraining by discarding the lowest-confidence queries ordered by the confidence of the corresponding token at the output of the encoder. As shown in Figure 4, this yields meaningful latency-accuracy tradeoffs. In addition, prior work (Zhao et al., 2024) demonstrates that decoder layers can be pruned at test time, since each layer is supervised independently during training. We find that it is possible to remove all decoder layers, relying solely on the initial query proposals from the two-stage DETR pipeline. In this case, there is no cross-attention to the encoder states or self-attention between queries, leading to a substantial runtime reduction. The resulting model resembles a single-stage YOLO-style architecture without NMS. Eliminating the final decoder layer reduces latency by 10% with only a 2 mAP drop in performance.
@@ Appendix C Benchmarking FLOPs
We benchmark FLOPs for RF-DETR, GroundingDINO, and LLMDet with PyTorch’s FlopCounterMode. We find that FlopCounterMode closely reproduces FLOP counts obtained with custom benchmarking tools for YOLOv11, D-FINE, and LW-DETR. In practice, we also find that it provides more reliable results than CalFLOPs (Ye, 2023). Notably, LW-DETR’s FLOPs count is roughly twice that of the originally reported result (cf. Table 9). We posit that this discrepancy can be attributed to LW-DETR reporting MACs instead of FLOPs. We rely on the officially reported FLOPs counts from YOLOv11, YOLOv8, D-FINE, and RT-DETR.
@@ Appendix D Impact of Class-Names on Open-Vocabulary Detectors
We evaluate the impact of fine-tuning open-vocabulary detectors like GroundingDINO with class names on RF100-VL in Table 10. Intuitively, GroundingDINO’s vision-language pre-training should be more more useful when we prompt with class names (e.g. car, truck, bus) instead of class indices (e.g. 0, 1, 2). Using class names when fine-tuning provides more information to the VLM about the underlying data than is available to non-VLM detectors, potentially leading to better downstream performance. However, we find that fine-tuning GroundingDINO on RF100-VL yields nearly identical performance in both cases, suggesting that naively fine-tuning the end-to-end model mitigates the benefits of open-vocabulary pre-training. Future work should investigate ways of effectively fine-tuning VLMs to preserve foundational pre-training.
@@
GroundingDINO (Liu et al., 2023) w/ Standard Class Names T 173.0M 1008.3 309.9 62.3 88.8 67.8 39.2 57.7 69.5
GroundingDINO (Liu et al., 2023) w/ Class Index Names T 173.0M 1008.3 309.9 62.5 88.2 68.3 40.0 58.4 70.3
@@ Appendix E Benchmarking Larger Model Variants
Detectors like LW-DETR (Chen et al., 2024a) and D-FINE (Peng et al., 2024) hand-design larger variants to scale up a model family. In contrast, NAS-based architectures like RF-DETR automatically discover scaling strategies through grid-based search. We analyze two families of RF-DETR models derived from distinct scaling strategies: one based on a DINOv2-S backbone and another based on a DINOv2-B backbone. To evaluate how well each family scales, we compare their NAS-generated Pareto curves against those of D-FINE. Specifically, at each D-FINE size, we identify the RF-DETR variant with the same backbone that maximizes performance at a comparable latency. For example, when comparing to D-FINE (small), we select the RF-DETR model that offers the best accuracy without exceeding D-FINE (small)’s latency. Note that these RF-DETRmodels are different than those reported in Tables 2 and 12.
As shown in Table 11, the DINOv2-S backbone family initially surpasses D-FINE in mAP@50:95 but fails to maintain this advantage at larger model sizes, suggesting that its scaling strategy is less effective than D-FINE’s manual design. In contrast, the DINOv2-B backbone family shows the opposite trend, where the performance gap between D-FINE and RF-DETR narrows as latency increases. This implies that at higher latencies, the DINOv2-B based RF-DETR models could surpass D-FINE (and indeed RF-DETR (2x-large) outperforms D-FINE on mAP 50:95). Importantly, expanding the D-FINE model family would require substantial additional engineering effort, whereas extending the RF-DETR model family is straightforward; higher-latency variants can be sampled directly from the same NAS search without re-training. We present the COCO and RF100-VL results of our larger variants in Tables 12, 13, and 14. We also include an RF-DETR Max variant on each dataset to show our method’s maximum performance with latency less than 100ms, a scale other model families don’t reach.
@@
YOLOv8 (Jocher et al., 2023) L 43.7M 165.2 8.0 49.5 64.7 54.0 30.2 55.1 68.5
YOLOv11 (Jocher et al., 2024) L 25.3M 86.9 6.5 49.9 64.9 54.5 30.4 55.9 68.1
YOLOv8 (Jocher et al., 2023) XL 68.2M 257.8 11.3 50.5 65.6 55.1 30.0 56.2 69.5
YOLOv11 (Jocher et al., 2024) XL 56.9M 194.9 10.5 50.9 66.1 55.4 31.5 56.6 68.7
@@
RT-DETR (Zhao et al., 2024) R50 42M 136 8.5 55.0 73.3 59.8 37.9 59.7 71.6
LW-DETR (Chen et al., 2024a) L 46.8M 137.5 6.9 56.1 74.6 61.0 37.1 60.4 73.0
D-FINE (Peng et al., 2024) L 31M 91 7.5 57.2 74.9 62.2 40.6 61.4 73.7
@@
RT-DETR (Zhao et al., 2024) R101 76M 259 12.0 56.1 74.5 61.1 38.1 60.4 73.4
LW-DETR (Chen et al., 2024a) XL 118.0M 342.5 13.0 58.3 76.9 63.3 40.2 63.3 74.7
D-FINE (Peng et al., 2024) XL 62M 202 11.5 59.3 76.8 64.6 42.1 64.2 76.3
@@
YOLOv8 (Jocher et al., 2023) L 46.0M 220.5 9.7 39.0 60.5 41.7 18.0 44.7 57.8
YOLOv11 (Jocher et al., 2024) L 27.6M 132.2 8.3 39.5 61.5 42.1 18.6 45.5 59.4
YOLOv8 (Jocher et al., 2023) XL 71.8M 344.1 14.0 39.5 61.3 42.1 18.9 45.6 58.8
YOLOv11 (Jocher et al., 2024) XL 62.1M 296.4 13.7 40.1 62.4 42.6 18.8 46.4 60.1
@@
YOLOv8 (Jocher et al., 2023) L 43.7M 165.2 7.9 56.5 82.1 61.1 7.1 46.0 48.9
YOLOv11 (Jocher et al., 2024) L 25.3M 86.9 6.4 56.5 82.2 61.0 6.4 45.5 49.0
YOLOv8 (Jocher et al., 2023) XL 68.2M 257.8 11.2 56.5 82.3 61.0 6.6 45.7 47.9
YOLOv11 (Jocher et al., 2024) XL 56.9M 194.9 10.3 56.2 81.7 60.8 6.1 45.9 48.1
@@
RT-DETR (Zhao et al., 2024) R50 42M 136 8.4 61.7 87.7 66.9 38.1 57.1 69.4
LW-DETR (Chen et al., 2024a) L 46.8M 137.5 6.8 61.5 87.4 67.0 37.1 56.4 69.0
D-FINE (Peng et al., 2024) L 31M 91 7.5 61.6 86.4 67.2 37.8 56.5 70.1
RF-DETR (Ours) L 34.1M 119.1 6.2 62.0 88.1 67.3 36.9 57.1 70.2
RF-DETR (Ours) w/ Fine-Tuning L 34.1M 119.1 6.2 62.3 88.2 67.4 37.1 57.2 70.3
RT-DETR (Zhao et al., 2024) R101 76M 259 11.9 61.0 87.4 66.2 36.6 56.3 68.2
LW-DETR (Chen et al., 2024a) XL 118.0M 342.5 13.0 62.1 87.9 67.6 37.4 57.1 70.2
D-FINE (Peng et al., 2024) XL 59.3 76.8 11.4 62.2 86.9 68.0 37.6 57.4 69.7
RF-DETR (Ours) XL 35.0M 199.0 9.7 62.6 88.5 67.9 39.0 57.8 70.4
RF-DETR (Ours) w/ Fine-Tuning XL 35.0M 199.0 9.7 63.0 88.7 68.2 38.8 58.2 70.6
@@
Appendix F Per-Knob Sensitivity Analysis
We evaluate the impact of varying resolution and patch size in Figure 5. Both curves follow a clear Pareto frontier, and are consistent with findings from prior work like FlexiViT (Beyer et al., 2023). RF-DETR is able to gracefully interpolate between seen (blue circles) and unseen (red stars) configurations during inference. Importantly, unseen configurations closely track the trend established by the seen configurations, demonstrating that RF-DETR generalizes beyond the model configurations encountered during training.
Figure 5: Per Knob Sensitivity Analysis. Despite never seeing certain resolutions and patch sizes, RF-DETR is able to gracefully interpolate to novel model configurations.
Appendix G Impact of NAS Fine-Tuning on COCO
@@
Appendix H Impact of Dataset Characteristics on Tunable Knobs
We evaluate the impact of different dataset characteristics on optimal network configurations for RF-DETR (medium) on RF100-VL in Table 17 and Figure 6. We compare different combinations of object size, number of spatial locations, number of decoder layers, number of windows, number of classes, number of annotations, objects per image, and number of queries below. We do not expect these relationships to be linear, but expect that they will be monotonic (e.g. non-zero slope). For example, we find strong correlations between the number of classes and number of decoder layers, objects per image and number of queries, spatial locations and number of windows. Notably, we do not find strong correlations between object size and number of decoder layers, number of annotations and number of decoder layers, and objects per image and number of decoder layers
Figure 6: Impact of Dataset Characteristics on Tunable Knobs. We visualize the correlation between key dataset characteristics and several tunable architectural knobs above. Across all subplots, individual points represent different datasets within RF100-VL while dashed lines show corresponding linear trends. Overall, the results indicate that object-centric properties of datasets such as average object size, number of classes, object density, and total annotations tend to have only modest influence on architectural choices like the number of decoder layers, number of windows, number of queries, and spatial resolution. Slight positive or negative trends appear in some cases (e.g., more classes or more objects per image loosely correlating with deeper decoders or higher query counts), but the scatter remains wide, suggesting no strong deterministic relationship. These findings highlight that while dataset characteristics offer some intuition for selecting model hyperparameters, optimal configurations ultimately depend on a combination of factors rather than any single dataset attribute.
Table 17: Regression Analysis. We evaluate the linear relationships between various dataset characteristics and model parameters. Although these correlations are non-linear, the line-of-best-fit helps explain general trends. P-values indicate the significance of the correlation.
Relationship Slope Intercept R-squared P-value Std. Error
Average Object Size vs Number of Spatial Locations -0.009 0.384 0.148 0.000 0.002
Average Object Size vs Number of Decoder Layers -0.000 0.044 0.000 0.971 0.006
Average Object Size vs Number of Windows -0.008 0.065 0.019 0.170 0.006
Number of Classes vs Number of Decoder Layers 0.837 2.125 0.026 0.106 0.513
Number of Annotations vs Number of Decoder Layers 698.763 6654.795 0.001 0.706 1848.733
Average Objects per Image vs Number Decoder Layers 0.163 7.618 0.000 0.857 0.904
Average Objects per Image vs Number of Queries 0.020 4.457 0.019 0.171 0.015
Number of Spatial Locations vs Number of Windows 1.722 32.984 0.492 0.000 0.177
Appendix I Impact of Fixed Architecture on RF100-VL
We evaluate the impact of transferring a NAS architecture optimized for COCO to RF100-VL in Table 18 and Figure 7. We find that these fixed architecture models perform remarkably well without further dataset-specific NAS. Specifically, RF-DETR (large) model with a fixed architecture achieves the best performance among all prior real-time models on COCO. However, dataset-specific NAS yields significant improvements. Notably, the performance delta from LW-DETR to the fixed architecture is comparable to the improvement from the fixed architecture to the NAS-optimized model on the target dataset for nano, small, and medium scale models.